### v1

@@ 5.2. Controlling explicit and personalized social media feeds
A central focus of Bonsai was to give participants greater control over their feeds. Our findings show that while participants valued this agency, they also grappled with the effort required to maintain it. This tension between effort and control emerged as a recurring theme: some participants found the work of curating feeds worthwhile for the increased precision it afforded, while others felt the costs outweighed the benefits once their feeds reached a “good enough” state.
@@
Participants also mentioned how much easier it is to use Bonsai compared with other feed building tools available for Bluesky, like graze.social or Skyfeed. For example P15, who has a lot of experience using feed building systems, said “it would have taken me probably hours in between creating, backfilling it, putting it in production, and then fine-tuning it to work. In here, in five minutes, you already have something that is a super good, almost final product.” Interestingly, the other feed building systems available used are rule-based; in contrast, using Bonsai was “super natural” and doesn’t require you to go into logic mode and start thinking about the building process by itself.”
@@
After the initial feed creation process, not all participants had the same motivation to keep editing their feed. Overall, the amount of control and enjoyment that participants felt when building and maintaining feeds with Bonsai seemed correlated to the amount of effort they subsequently invested. On one side, four participants chose to put in significant effort to try to improve their feeds and reported that this effort was worth the improvements. P11, for example, took “an extra 5 minutes” to manually add a subset of accounts to the list of sources after the initial suggestions, and describes her experience:
@@
Five participants were satisfied with their initial feed, made minimal edits, and then estimated that additional effort was not worth improvement in their feeds. P6 is one such lightweight user. Once they found their feed useful enough and didn’t encounter any “deeply upsetting” content, they did not feel the need to keep tweaking it, they said “I think the effort to reward or change ratio no longer made sense, […] I wouldn’t go in to be like, ‘okay, I don’t want job posts’.” P8 enjoyed that with the natural language interface “if I can just put text, it’s like, okay, I [wrote] two sentences, or let it be 10 sentences, and then I have, kind of my preferences captured.” Yet, P8 felt that further fine-grained edits were unnecessary, saying “having too much individual control is too much effort, […] I don’t think I would want to put that much effort into engineering my feed.”
@@
Finally, for the six remaining participants, their initial feed did not meet their expectations, and they ultimately either stopped using it or created a new feed where they had more success. This was the case for P10, who prompted the system to suggest “new accounts that are followed by people [they] currently follow” and asked for “more users local to [them]”. However, the current system is not equipped with account suggestions based on existing accounts and lacks data on the location of accounts or posts. Due to these restrictions, P10 built a final feed centered around a niche topic of print media and zines and found this feed “more enjoyable.” They explained that their first feed “felt overwhelming a bit, because everything seems totally new from a person that I didn’t know” and that in its current state they would use the system “if [they] want to look for something totally new […] [that they] don’t have as much of in [their] current account”. Another example is P3, who tried recreating a feed similar to their ‘following’ feed but noticed that “the content that I engage most with from this feed is coming from the accounts I already followed”, which led them to prioritize continuing to use their ’following’ feed instead of their Bonsai feed. P3 explained that they felt the effort trade-off was not worth it with the time they had: “it’s a hard trade-off, because time is both what I gain, but also what I would need to put in to make it better.”
@@
Despite the effort required to create feeds that fully meet user intentions, some participants in our study felt empowered and excited by the ability to curate their feed. P4 said “it is empowering to actually have a different tool in the toolbox, which you can use to create, sort of, your own… Custom universe, if you will.” Similarly, for P11 “this was a true delight and joy to use. The college sports [feed] will be my default feed if I can continue to access this through for the next 6 months—truly.” However, satisfaction with Bonsai did not prevent participants from feeling that the cognitive load of having to formulate their intentions was too high. For example P13 was “very happy with the ‘limit posts about’ when it comes to political content and news.” Reflecting about his science-related news feed, they said “I would only see science-related news, and that made me very happy.” Yet, they also expressed frustration that the feeds could not learn based on their on-platform behaviors, saying “I’m very lazy and I want to be more in control, but I also want to continue to be lazy in my social media experience.”
@@
5.3. Procedural transparency of explicit and personalized feeds
@@ 5.3.2. Participants questioned how natural language instructions would be translated to feed outcomes
Others expressed doubt about whether language models could handle complex or ambiguous requests. P8 felt “a bit uncomfortable giving too many instructions” and was “not sure whether current language models are actually that good at being able to distinguish that,” particularly when combining diverse topics. The participant questioned whether the system could reliably “decide what is interesting and what is not.” For P1, who worked with NLP, these concerns were even more pronounced:
@@
These reflections highlight how assumptions about natural language processing can create hesitation and self-consciousness when formulating prompts, especially for technically knowledgeable users. At the same time, participants noted that this challenge might be less pronounced for more general audiences. For example, P1 also noted that ”Depending on the intended audience, that [previous assumptions] might not be as big a point of friction for other people.”
@@ 6. Discussion
By examining Bonsai, we explore how intentional and personalized feedbuilding offers an alternative vision for the future of social media. Our findings show that participants used intentional feeds both to filter out toxic or irrelevant content and to surface new, meaningful material that was otherwise hard to find. Participants also reported feeling relieved from the pressure of having their engagement shape future content, but also noted that building feeds required more effort than they were accustomed to. Taken together, these results outline a design space for intentional social media feeds, including fundamental tradeoffs that arise from shifting fine-grained control more directly to users. Furthermore, they extend prior critiques of engagement-optimized feeds that document addictive or regrettable use (Tran et al., 2019a; Cho et al., 2021), demonstrating how explicit, intention-first personalization can offer benefits.
@@
This work advances an intention-first paradigm for creating and engaging with social media feeds. In this paradigm, systems construct feeds transparently around user goals defined directly instead of inferring preferences from engagement. Our findings show that participants often used Bonsai in two complementary ways: to filter out toxic or irrelevant content that undermined their well-being, and to broaden their horizons by surfacing new material that aligns with their interests. This dual role distinguishes intentional feeds from engagement-based systems that can encourage compulsive scrolling and erode well-being (Cho et al., 2021; Tran et al., 2019b). In contrast to prior interventions, which primarily sought to reduce screen time (Hiniker et al., 2016; Ko et al., 2015; Grüning et al., 2023), we show how intentional feeds can redirect attention toward exploration, diversification, and discovery—not just restraint.
@@
Our results indicate that participants felt liberated from the burden of knowing their behavioral actions (likes, comments, replies) could shape future feed content. Prior work shows that people develop “folk theories” to reason about algorithmic curation (DeVito et al., 2017; Gelman and Legare, 2011; Karizat et al., 2021). Our study highlights that these theories themselves can impose cognitive load, as users monitor how each click might ”teach” the system and constrain their interactions. By decoupling intentions from engagement, Bonsai relieved this burden, enabling participants to treat liking, replying, or reposting as expressive rather than instrumental. Our findings thus suggest that transparency and explicit intent-setting do more than improve comprehension; they can also reshape the affective experience of social media use by restoring freedom and spontaneity to online interactions, a shift that resonates with broader calls for more agentic, trust-building forms of transparency and control (Ananny and Crawford, 2018).
@@
At the same time, our study highlights that intentional feeds carry their own cognitive costs, which may be unavoidable rather than incidental. Participants often described the act of articulating preferences as effortful, echoing findings from the personal informatics literature where users weigh the burden of self-tracking against the value of self-improvement (Epstein et al., 2015). This finding resonates with broader models of technology adoption, which emphasize the role of effort expectancy in shaping use (Camilleri, 2024). With Bonsai, natural language input substantially lowered the barrier to entry, but participants still wrestled with how best to phrase instructions or interpret the system’s responses. Here, procedural transparency plays a critical role: clearer visibility into how inputs map to outputs can reduce uncertainty, reinforce users’ belief in the system’s efficacy, and make the effort of intentionality feel more worthwhile. Yet our findings underscore a more fundamental tradeoff: emergent feeds, which dominate today’s platforms, offer immediacy and ease but at the expense of control (Cunningham et al., 2024; Bandy and Lazovich, 2023); explicit feeds provide agency and alignment but demand greater effort, as also observed in prior systems that required active curation (Bernstein et al., 2010; He et al., 2023; Davis et al., 2023). Future systems could grapple with this tension directly, considering hybrid approaches that combine the ease of emergent feeds with the fidelity of intentional ones.
@@ 7. Conclusion
8. Generative AI Disclosure
@@ Appendix A Additional User Interface Screenshots
We present additional user interface screenshots in Figure 7.
Appendix B Source Database Construction

### v2

@@ 5.2. Steering social media feeds
A central focus of Bonsai was to give participants a simple way to construct personalized feeds. Our findings show that while participants valued this agency, they also grappled with the effort required to maintain it. This tension between effort and reward emerged as a recurring theme: some participants found the work of curating feeds worthwhile for the increased precision it afforded, while others felt the costs outweighed the benefits once their feeds reached a “good enough” state. Participants with prior feedbuilding experience found Bonsai comparably easy to use, while those used to emergent feeds could struggle with the magnitude of the paradigm shift.
@@
Participants with experience with other feedbuilders specifically mentioned how much easier it is to use Bonsai compared with other feed building tools available for Bluesky, like graze.social or Skyfeed. P6 enjoyed that Bonsai easily facilitated setting up a feed, and observed that setting up custom feeds through Skyfeed or graze.social seemed like “a lot up front to not even know if you’re gonna get a particularly good feed out of it.” P15 has a lot of experience using feed building systems, and said “it would have taken me probably hours in between creating, backfilling it, putting it in production, and then fine-tuning it to work. In here, in five minutes, you already have something that is a super good, almost final product.” Specifically, the other feed building systems available to use are all rule-based. P11, who had used Skyfeed in the past, explained that in other interfaces they were “just basically writing Boolean operators about these combinations of words,” whereas Bonsai let them “just kind of stream of consciousness talk about the things you want to see, which is much more intuitive, right? It gets you closer to that— I know what I want to see, rather than trying to sit there and think of keywords.” Similarly, P15 described using Bonsai as “super natural:” “doesn’t require you to go into logic mode and start thinking about the building process by itself.” In all, compared to current state-of-the-art feed builders, participants described Bonsai as easier to use, particularly at setup.
@@
After the initial feed creation process, not all participants had the same motivation to keep editing their feed. Overall, the amount of control and enjoyment that participants felt when building and maintaining feeds with Bonsai seemed correlated to the amount of effort they subsequently invested. On one side, four participants chose to put in significant effort to try to improve their feeds and reported that this effort was worth the improvements. P11, for example, took “an extra 5 minutes” to manually add a subset of accounts to the list of sources after the initial suggestions, and describes their experience:
@@
Five participants were satisfied with their initial feed, made minimal edits, and then estimated that additional effort was not worth improvement in their feeds. P6 is one such lightweight user. Once they found their feed useful enough and hadn’t encountered any “deeply upsetting” content, they did not feel the need to keep tweaking it, saying “I think the effort to reward or change ratio no longer made sense, […] I wouldn’t go in to be like, ‘okay, I don’t want job posts’.” P8 enjoyed that with the natural language interface “if I can just put text, it’s like, okay, I [wrote] two sentences, or let it be 10 sentences, and then I have, kind of my preferences captured.” Yet, P8 felt that further fine-grained edits were unnecessary, saying “having too much individual control is too much effort, […] I don’t think I would want to put that much effort into engineering my feed.”
@@
Finally, for the six remaining participants, their initial feed did not meet their expectations, and they ultimately either stopped using it or created a new feed where they had more success. This was the case for P10, who prompted the system to suggest “new accounts that are followed by people [they] currently follow” and asked for “more users local to [them]”. However, the current system is not equipped with account suggestions based on existing accounts and lacks data on the location of accounts or posts. While the system still tried to suggest relevant content to the user, future iterations should more clearly communicate system constraints. Due to these restrictions, P10 built a final feed centered around a niche topic of print media and zines and found this feed “more enjoyable.” They explained that their first feed “felt overwhelming a bit, because everything seems totally new from a person that I didn’t know.” Another example is P3, who tried recreating a feed similar to their ‘following’ feed but noticed that “the content that I engage most with from this feed is coming from the accounts I already followed”, which led them to prioritize continuing to use their ’following’ feed instead of their Bonsai feed. P3 explained that they felt the effort trade-off was not worth it with the time they had: “it’s a hard trade-off, because time is both what I gain, but also what I would need to put in to make it better.”
Participants thus had three different types of experiences, and generally achieved more rewarding feeds when they put more effort into learning and updating the Bonsai system. We illustrate these divergent types of user journeys in Figure 5. Though we cannot draw any definite quantitative conclusions, our interviews suggest that the two reasons that motivated participants’ willingness to invest additional effort were prior experience with custom feed builders and a strong desire for specific feeds. P13 had never used other feed builders, but was dissatisfied with Bluesky’s default ‘Discover’ feed, and was therefore highly motivated to create a new feed, finding that their Bonsai feed is “way better than the Discover feed on Blue Sky.” Prior experiences with feed-building tools also separated participants who put in extra time to refine their feeds from participants who did not. As we highlight in Section 5.2.1, those who were used to other feedbuilders were accustomed to effortful feed creation experiences, and appreciated the comparative flexibility and ease of using Bonsai. In contrast, participants with limited experience with other feed builders brought only their expectations from emergent feeds, which require minimal effort. For these participants, spending time constructing an intentional feed represented a paradigm shift that did not align with their expectations. P10 captured how intentional feed editing can seem quite intimidating for novice users: “it feels like there’s a bit of a learning curve for how to use this tool, because […] we’re expecting to be fast when we use [feeds].”
Figure 5. Three representative participant user journeys using Bonsai: low effort, low reward (top row); low effort, medium reward (middle row), high effort, high reward (bottom row). The top two participants (P10 and P7) represent perspectives of people who spent minimal effort to improve their feeds. P11 (bottom) represents a participant who had prior experience using other feed builders, and invested time to optimize their Bonsai feeds and reap high rewards.
A diagram showing three user scenarios for feed customization, each with Intention, Effort, and Outcome columns. Participant 10 (orange row) wanted a feed based on their current network and nearby people, exerted low effort, and experienced low reward, feeling overwhelmed and valuing social connections over content. Participant 7 (gray row) wanted deep analysis on experimental or AI films, exerted low effort, and achieved medium reward, finding 30% very interesting, 30% somewhat interesting, and 40% uninteresting. Participant 11 (teal row) wanted a feed for only their college basketball teams, exerted high effort by iterating on three different feeds, and achieved high reward, enjoying the final result.
@@
Despite the effort required to create feeds that fully meet user intentions, some participants in our study felt empowered and excited by the ability to curate their feed. P4 said “it is empowering to actually have a different tool in the toolbox, which you can use to create, sort of, your own… Custom universe, if you will.” Similarly, for P11 “this was a true delight and joy to use. The college sports [feed] will be my default feed if I can continue to access this through for the next 6 months—truly.” However, satisfaction with Bonsai did not prevent participants from feeling that the cognitive load of having to formulate their intentions was too high. For example, P13 was “very happy with the ‘limit posts about’ when it comes to political content and news.” Reflecting on their science-related news feed, they said “I would only see science-related news, and that made me very happy.” Yet, they also expressed frustration that the feeds could not learn based on their on-platform behaviors, saying “I’m very lazy and I want to be more in control, but I also want to continue to be lazy in my social media experience.”
@@
5.3. Procedural transparency of intentional and personalized feeds
@@ 5.3.2. Participants questioned how natural language instructions would be translated to feed outcomes
Others expressed doubt about whether language models could handle complex or ambiguous requests. P8 felt “a bit uncomfortable giving too many instructions” and was “not sure whether current language models are actually that good at being able to distinguish that,” particularly when combining diverse topics. The participant questioned whether the system could reliably “decide what is interesting and what is not.” For P1, who worked in natural language processing, these concerns were even more pronounced:
@@
These reflections highlight how assumptions about natural language processing can create hesitation and self-consciousness when formulating prompts, especially for technically knowledgeable users. At the same time, participants noted that this challenge might be less pronounced for more general audiences. For example, P1 also noted that “depending on the intended audience, that [previous assumptions] might not be as big a point of friction for other people.”
@@ 6. Discussion
By examining Bonsai, we explore how intentional and personalized feedbuilding offers an alternative vision for the future of social media. Our findings show that participants used intentional feeds both to filter out toxic or irrelevant content and to surface new, meaningful material that was otherwise hard to find. Participants also reported feeling relieved from the pressure of having their engagement shape future content, but noted that building feeds required more effort than they were accustomed to. Taken together, these results outline a design space for intentional social media feeds, including fundamental tradeoffs that arise from shifting fine-grained control more directly to users. Furthermore, they extend prior critiques of engagement-optimized feeds that document addictive or regrettable use (Tran et al., 2019b; Cho et al., 2021), demonstrating how explicit, intention-first personalization can offer benefits.
6.1. Promoting Intentional Goal Alignment
@@
This work advances an intention-first paradigm for creating and engaging with social media feeds. In this paradigm, systems construct feeds transparently around user goals defined directly instead of inferring preferences from engagement. Our findings show that participants often used Bonsai in two complementary ways: to filter out toxic or irrelevant content that undermined their well-being, and to broaden their horizons by surfacing new material that aligns with their interests. This dual role distinguishes intentional feeds from engagement-based systems that can encourage compulsive scrolling and erode well-being (Cho et al., 2021; Tran et al., 2019a). In contrast to prior interventions, which primarily sought to reduce screen time (Hiniker et al., 2016; Ko et al., 2015; Grüning et al., 2023), we show how intentional feeds can redirect attention toward exploration, diversification, and discovery, rather than restraint. We find that Bonsai’s natural language interface encouraged such playfulness and nuanced preference specifications, in ways that other custom feed builders currently lack.
@@
Our results indicate that participants felt liberated from the burden of knowing their online engagement (likes, comments, replies) could shape future feed content. Prior work shows that people develop “folk theories” to reason about algorithmic curation (DeVito et al., 2017; Karizat et al., 2021). Our study highlights that these theories themselves can impose cognitive load, as users monitor how each click might ”teach” the system and constrain their interactions. By decoupling intentions from engagement, Bonsai relieved this burden, enabling participants to treat liking, replying, or reposting as expressive rather than instrumental. Our findings thus suggest that transparency and explicit intent-setting on social media feeds can help users reap more pleasure from their online experiences by encouraging freedom and spontaneity. Today, half of U.S. teens believe that social media is detrimental to people their age: our research shows one path to making social media fun again. (Faverio et al., 2025) Our work thus resonates with broader calls for more agentic, trust-building forms of transparency and control (Ananny and Crawford, 2018).
6.2. Effort-Reward Tradeoffs
@@
At the same time, our study highlights that intentional feeds carry their own cognitive costs, which may be unavoidable rather than incidental. Participants often described the act of articulating preferences as effortful, echoing findings from the personal informatics literature where users weigh the burden of self-tracking against the value of self-improvement (Epstein et al., 2015). We identify that participants willing to overcome this barrier largely had strong motivations or more pragmatic expectations of effort based on prior experiences with feed builders. With Bonsai, natural language input substantially lowered the barrier to entry compared with other popular feed builders, but participants still wrestled with how best to phrase instructions or interpret the system’s responses. Here, procedural transparency plays a critical role, since clearer visibility into how inputs map to outputs can reduce uncertainty, reinforce users’ belief in the system’s efficacy, and make the effort of intentionality feel more worthwhile. Our findings directly echo Vroom (1964)’s expectancy theory of motivation, which emphasizes the role of effort expectancy, instrumentation, and reward expectancy to explain what drives people to take effortful actions. Yet, our findings also underscore a potentially fundamental tradeoff: emergent feeds, which dominate today’s platforms, offer immediacy and ease but at the expense of control (Cunningham et al., 2024; Bandy and Lazovich, 2023); intentional feeds provide agency and alignment but demand greater effort, as also observed in prior systems that required active curation (Bernstein et al., 2010; He et al., 2023; Davis et al., 2023). Future systems could also consider hybrid approaches that combine the ease of emergent feeds with the fidelity of intentional ones.
6.3. The Future of Intentional Feeds
Intentional social media feeds will be most impactful if they are broadly and meaningfully adopted by technology companies and individuals. This work emerged from the belief that the current engagement-first paradigm is more aligned with business incentives than human values. However, Kleinberg et al. (2024) argue that there are real financial costs to over-optimizing for engagement: users may quit the platform, thereby reducing all future earnings potential. Additionally, technology platforms increasingly face public and legal pressure to demonstrate how they prevent harmful mental health impacts (Commonwealth of Massachusetts, 2023; Ball, 2025). Whether motivated by financial outcomes or public pressure, the largest technology companies have begun to tease features that better align social feeds with users’ long-term wellbeing and goals. Since the writing of this paper, Meta’s Threads and Instagram have announced features to give users more granular control over the topics they see in their feed (Perez, 2025; Silberling, 2025), while X’s CEO Elon Musk promised that users will be able to “adjust [their] feed dynamically just by asking Grok” (Musk, 2025). One plausible and even likely future for intentional feeds is therefore that they become increasingly incorporated into existing social media products. In such a future, one risk would be that these tools become a form of “user agency theatre” (features that lend the impression of increasing user agency but in practice have a limited impact on the outcome of interest (Schneier, 2003)), where companies use the presence of seldom-used features to maintain their current profitable and algorithmic approaches.
At the same time, the architecture we deploy in Bonsai is unlikely to be directly adopted at platform scale. Running large LMs across planning, sourcing, and per-post curation is both computationally expensive and latency-sensitive, especially for real-time, high-throughput feeds. In practice, any production deployment inspired by Bonsai would need to dramatically reduce LM calls—for instance, by using embeddings and retrieval-style architectures to pre-compute user and item representations. In this sense, Bonsai should be understood as a pilot that probes what intent-first, LM-mediated feeds can feel like for users, and that surfaces design and engineering directions for hybrid systems in which LMs help set and refine user goals while cheaper models handle most of the online ranking workload.
@@ 6.5. Limitations
Fifth, Bonsai was built and optimized for content-based filtering. Users sometimes found new use cases we had not anticipated, such as asking Bonsai to fetch and analyze their existing network. In such unanticipated scenarios, system restrictions could cause friction for users, who may not have known their request could not be translated to a feed outcome. This system constraint partly motivates this paper’s emphasis on procedural transparency.
@@ 7. Conclusion
Acknowledgements.
We thank the Hasler Foundation for partial support of OEM during his master’s thesis, and the Cornell Tech Social Technologies Lab for feedback on earlier versions of this work.
@@ Appendix A Additional User Interface Screenshots
We show additional screenshots of Bonsai’s user interface on Figure  6 and Figure  7.
Appendix B Database construction
@@
(a) Feed configuration page
(b) Manage Feeds page
Figure 7. Ranking presets in Bonsai. Users can choose between four styles: Balanced (equal emphasis on relevance, popularity, and recency), Focused (prioritizes relevance), Fresh (emphasizes recency), and Trending (emphasizes popularity). Each preset adjusts point weights that contribute to the final ranking algorithm.
This figure shows the four ranking presets that Bonsai offers to let users control how their feeds are ordered. Each preset adjusts the relative weights assigned to three criteria: relevance, popularity, and recency. The “Balanced Feed” splits weights fairly evenly, with a slight emphasis on relevance, producing an even mix of relevant, popular, and recent content. The “Focused Feed” heavily favors relevance, giving less weight to popularity and recency, so that content closely aligned with the user’s stated goals is prioritized. The “Fresh Feed” emphasizes recency, surfacing the newest posts while still considering relevance and popularity to a smaller degree. Finally, the “Trending Feed” highlights popularity, prioritizing posts that are widely engaged with, while still incorporating recency and relevance. Each point in the weight system contributes incrementally to the ranking formula, allowing these presets to serve as simple yet powerful shortcuts for tailoring how Bonsai assembles a user’s feed. This figure presents two core user interfaces from Bonsai that support feed customization and management. On the left, the feed configuration page illustrates how users create a custom feed, here named “Paw Paradise.” The interface allows users to select sources such as accounts, hashtags, or search terms, specify inclusion rules like always showing posts with cats and dogs, and set exclusion rules such as hiding distressing or promotional content. At the bottom, users can also choose a ranking style (Focused, Fresh, Balanced, or Trending) depending on whether they want to prioritize relevance, recency, or popularity. On the right, the Manage Feeds page shows how users can oversee multiple feeds at once, with examples like ”Professional Following” and ”Paws & Whiskers.” From this page, users can edit, activate, duplicate, or delete feeds, and view their active feed directly on Bluesky. Together, these pages demonstrate how Bonsai combines fine-grained configuration with simple management tools for intentional, personalized feed building.